compound is found to contain 63.65 % nitrogen and 36.35 % oxygen by mass. the empirical formula is? the molar mass is 44.02 g/mol. what is the molecular formula?

A compound is found to contain 6.360 % silicon , 36.18 % bromine , and 57.46 % iodine by mass. the empirical formula is? the molar mass is 441.7. what is the molecular formula?

A compound is found to contain 64.80 % carbon , 13.62 % hydrogen , and 21.58 % oxygen by mass. the empirical formula is? the molar mass is 74.14 g/mol, what is the molecular formula?
Answer #1

1)

we have mass of each elements as:

N: 63.65 g

O: 36.35 g

Divide by molar mass to get number of moles of each:

N: 63.65/14.01 = 4.543

O: 36.35/16.0 = 2.272

Divide by smallest to get simplest whole number ratio:

N: 4.543/2.272 = 2

O: 2.272/2.272 = 1

So empirical formula is:N2O

Molar mass of N2O,

MM = 2*MM(N) + 1*MM(O)

= 2*14.01 + 1*16.0

= 44.02 g/mol

Now we have:

Molar mass = 44.02 g/mol

Empirical formula mass = 44.02 g/mol

Multiplying factor = molar mass / empirical formula mass

= 44.02/44.02

= 1

So molecular formula is:N2O

Answer: N2O

2)

we have mass of each elements as:

Si: 6.36 g

Br: 36.18 g

I: 57.46 g

Divide by molar mass to get number of moles of each:

Si: 6.36/28.09 = 0.2264

Br: 36.18/79.9 = 0.4528

I: 57.46/126.9 = 0.4528

Divide by smallest to get simplest whole number ratio:

Si: 0.2264/0.2264 = 1

Br: 0.4528/0.2264 = 2

I: 0.4528/0.2264 = 2

  

So empirical formula is:SiBr2I2

Molar mass of SiBr2I2,

MM = 1*MM(Si) + 2*MM(Br) + 2*MM(I)

= 1*28.09 + 2*79.9 + 2*126.9

= 441.69 g/mol

Now we have:

Molar mass = 441.7 g/mol

Empirical formula mass = 441.69 g/mol

Multiplying factor = molar mass / empirical formula mass

= 441.7/441.69

= 1

So molecular formula is:SiBr2I2

Answer: SiBr2I2

3)

we have mass of each elements as:

C: 64.8 g

H: 13.62 g

O: 21.58 g

Divide by molar mass to get number of moles of each:

C: 64.8/12.01 = 5.3955

H: 13.62/1.008 = 13.5119

O: 21.58/16.0 = 1.3487

Divide by smallest to get simplest whole number ratio:

C: 5.3955/1.3487 = 4

H: 13.5119/1.3487 = 10

O: 1.3487/1.3487 = 1

  

So empirical formula is:C4H10O

Molar mass of C4H10O,

MM = 4*MM(C) + 10*MM(H) + 1*MM(O)

= 4*12.01 + 10*1.008 + 1*16.0

= 74.12 g/mol

Now we have:

Molar mass = 74.14 g/mol

Empirical formula mass = 74.12 g/mol

Multiplying factor = molar mass / empirical formula mass

= 74.14/74.12

= 1

So molecular formula is:C4H10O

Answer: C4H10O